14x-2y=46
-7x+y=-23

Answer by ikleyn(52776)

.
14x - 2y = 46
-7x + y = -23

These two equations are equivalent to one equation
-7x + y = 23.

And they have infinitely many solutions y = 23+7x for any (arbitrary) value of x.